# Superconductivity in Electron Liquids: Precision Many-Body Treatment of Coulomb Interaction

### v1

@@ VI Conventional Superconductors
We apply this framework to revisit e-ph SC in simple metals where properties of conduction electrons can be approximated by UEG. Existing predictions of $T_{\text{c}}$ based on the downfolded ME theory or derived approximations, such as the McMillan or Allen-Dynes formulas, result in order-of-magnitude uncertainties for simple metals with sub-Kelvin $T_{\text{c}}$, depending on the choice of the phenomenological Coulomb pseudopotential [morel_anderson].
@@ VI.1 Methods
Li(9R) 40 242444Electron-phonon coupling in lithium for 9R and hcp structures [liu_structural_1999]. We computed the hcp case, but also adapted results from literature for the 9R case. 0.34444Electron-phonon coupling in lithium for 9R and hcp structures [liu_structural_1999]. We computed the hcp case, but also adapted results from literature for the 9R case. 0.34444Electron-phonon coupling in lithium for 9R and hcp structures [liu_structural_1999]. We computed the hcp case, but also adapted results from literature for the 9R case. 3.25555Data for $r_{\text{s}}$ and $m_{\text{b}}$ in lithium were adapted from Ref.[richardson_effective_1997]. 1.75555Data for $r_{\text{s}}$ and $m_{\text{b}}$ in lithium were adapted from Ref.[richardson_effective_1997]. 0.18 $5\times 10^{-3}$ $4\times 10^{-4}$ 0.35
Li(hcp) 41 243 0.33444Electron-phonon coupling in lithium for 9R and hcp structures [liu_structural_1999]. We computed the hcp case, but also adapted results from literature for the 9R case. 0.37 3.19 1.4 0.17 0.03 $4\times 10^{-4}$ 0.64
@@
To obtain band structures and phonon spectra we employed Quantum Espresso (QE) [QE1, QE2, QE3] package. In all calculations the optimized norm-conserving Vanderbilt (ONCV) potential [ONCV1, ONCV2] and Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [perdew_generalized_1996] were used. Metallic crystals considered in this work form either FCC, or BCC, or HCP Bravais lattices. For the self-consistent-field (SCF) calculation, the energy cutoff was set to 90 Ry, the momentum-grid had $\times 24\times 24, 24×24$$\times 24\times 24$, $24\times 24\times 12$ points for the FCC, BCC, and HCP structures, respectfully. The error on the converged total energy was less than 0.001 Ry per atom. Phonons were calculated using the DFPT approach of Ref. [dfpt_rev], which was implemented within the QE package. A coarse $\mathbf{q}$-grid for the FCC, BCC, and HCP lattices contained $\times 6\times 6, 6×6$$\times 6\times 6$, and $6\times 6\times 3$ points, respectively, and the acoustic sum rule was used to eliminate the small imaginary frequency at the $\mathbf{\Gamma}$ point.
To get the e-ph coupling constant $\lambda$, we used the EPW package [EPW1, EPW2], which evaluates e-ph interaction for Wannier functions, and subsequent Fourier interpolation to obtain the e-ph interaction defined on an arbitrary $\mathbf{k},\mathbf{q}$ grid. Wannier functions were generated by the Wannier90 package [Wann1, Wannier_Vanderbilt_1997, Wannier_Vanderbilt_2001] using coarse $\mathbf{k}$-grids for the FCC, BCC, and HCP lattices with $\times 12\times 12, 12×12$$\times 12\times 12$, and $12\times 12\times 6$ points, respectively. The Wannier projectors for Li, Na, K, Mg, Al were set to $s$ and $p$ orbitals, and the ones for Zn are $s$, $p$, $d$ orbitals. The mismatch between the band structures interpolated from Wannier functions and QE was less than 0.01 eV in the energy window below (above) 1 eV around the Fermi level. To get a converged result for $\lambda$, we set a fine $\mathbf{k}$ ($\mathbf{q}$) grid for the FCC, BCC, and HCP lattices with $\times 60\times 60, 60×60$$\times 60\times 60$, and $60\times 60\times 30$ points, respectively, to have energy resolution at the level of 0.02 meV.
@@ VI.2 Results
Figure 10: The pressure dependence of the superconducting critical temperature in aluminum. The squares are our theoretical results; the lines are guides to the eye. Experimental data from Levy et al. [levy_can_1964] and Gubser et al. [gubser_high-pressure_1975] are plotted as diamond and circular markers, respectively.
In Fig. 10, we compare the experimental data for the pressure-dependent $T_{\text{c}}$ in aluminum [levy_can_1964, gubser_high-pressure_1975] with the corresponding results of our approach (up to the highest reported pressure). We see that our approach accurately captures the experimental trend, showing a clear decrease in $T_{\text{c}}$ as pressure increases from ambient conditions to $6$ GPa.
Although similar benchmarks using the SCDFT method [profeta_superconductivity_2006] also show good agreement with the experiment, the treatment of Coulomb repulsion within SCDFT involves uncontrolled approximations with respect to vertex corrections, screening dynamics, and quasiparticle weight. In contrast, our approach is based on a rigorous microscopic foundation and clear understanding of how one should proceed in cases significantly different from UEG. We performed calculations beyond the current experimental pressure limit of 66GPa; see Fig. 11. Assuming no structural transitions, we predict that SC in Al vanishes at pressure 60~60GPa.
@@
With the same numeric protocol, we calculated $T_{c}$ for other simple metals; the results are presented in Fig. 11 and TABLE 2. Given $T_{c}/\omega_{\text{log}}\propto\exp(-1/g)$ scaling, we choose to plot all data using $y=1/\ln(\omega_{\text{log}}/T\text{c})$ for the vertical axis, which may be considered as an effective attractive coupling in the Cooper channel (if $T_{\text{c}}$ is finite) for a given $\mu^{*}$ and $\lambda$ (here $\omega_{\text{log}}$ is the $\log$-averaged phonon frequency calculated following Ref. [allen1975]).
@@
Our results suggest that Mg, Na, and K are close to the critical point between states with and without s-wave SC. (We do not use the term “quantum critical point” here because SC in some other symmetry channel will preempt it.) Although these metals remain in the normal state as the temperature is lowered (e.g., well below $1K$), they enter a quantum critical regime characterized by a diverging pair-field susceptibility $\chi\sim\ln(T)$ [pengcheng] without any fine-tuning. This signal is worth verifying in future experiments.
@@ VII Conclusion
We developed a rigorous ab initio framework, Fig. 9, to obtain downfolded ME theory on the Fermi surface for correlated electrons, which resolves long-standing ambiguities in the treatment of Coulomb repulsion. By systematically integrating out high-energy electronic degrees of freedom using field-theoretical renormalization techniques, we arrive at frequency-only ME equations with all parameters precisely linked to the underlying microscopic Hamiltonian: the Coulomb pseudopotential $\mu^{*}$, determined by the two-quasiparticle scattering amplitude renormalized by the Bogoliubov-Tolmachev-Shirkov logarithm [bogoliubov59, tolmachev1961logarithmic], the electron-phonon coupling $\lambda$ screened by the quasiparticle vertex correction, and the quasiparticle density of states. Our approach eliminates the need for phenomenological—and thus uncontrolled—treatment of key parameters and establishes protocols for computing $\mu^{*}$ and $\lambda$ from the first principles.
@@
The ab initio framework of Fig. 9 applies not only to simple metals but to any other correlated material with well-defined quasiparticles at the Fermi surface and large separation between the phonon and electron energy scales $\omega_{\text{D}}\ll E_{\text{F}},\omega_{\text{p}}$. However, implementing high-order diagrammatics and renormalization, as it was done for UEG, is a major technical challenge for lattice systems with complex band structure lacking spherical symmetry. For example, in transition metals more than one band is involved in screening [schrieffer_1969] and Umklapp processes cannot be neglected. In cases including two dimensional systems the particle-particle channel Coulomb interaction is non-local (not well approximated by a structureless interaction and as noted by Simonato, Katsnelson and Rosner [Simonato23] a more careful treatment is required. If the energy-scale separation condition, on which the downfolding approximation rests, is not satisfied, e.g. at extreme densities relevant to astrophysical objects such as white and black dwarf stars [ginzburg_superconductivity_1968], or perhaps in flat-band systems and 2D materials where the $\sqrt{q}$ plasmon dispersion creates challenges to the energy scale separation arguments there may be no real alternative to working with full BSE.
Our findings highlight the crucial role of e-e correlations in the two-particle vertex function; they cannot be quantified by the single-electron exchange-correlation potential within the DFT approximation [kohn1964, kohn1965]. Modern numerical methods like VDiagMC enable accurate computation of these correlations, providing key insights into the topic of paring despite strong repulsive forces, and opening new possibilities for understanding other emergent phenomena, such as accurate calculation of many-body forces in metals, with direct implications for more reliable molecular dynamics simulations. In conclusion, by providing a complete and predictive first-principles framework for incorporating e-e correlations into the existing machinery of predicting superconductivity from e-ph interaction, this work represents a step forward in our ability to predict superconductivity starting from properties of the microscopic Hamiltonian without uncontrolled approximations.
@@ Acknowledgements.
We thank S. Das Sarma for valuable discussions. K. Chen and X. Cai are supported by the National Natural Science Foundation of China under Grants No. 12474245 and No. 12447103, the National Key Research and Development Program of China under Grant No.2024YFA1408604, and the GHfund A (202407010637). X. Cai and T. Wang acknowledge support from the National Science Foundation under Grant No. DMR-2335904. T. Zhang and S. Zhang acknowledge the support from National Key R&D Project (Grant Nos. 2023YFA1407400 and 2024YFA1409200), and the National Natural Science Foundation of China (Grant Nos. 12374165 and 12447101). A. Millis’s work at Columbia University was supported by the Keele Foundation and the Columbia University Materials Research Science and Engineering Center (MRSEC) through NSF Grant DMR-2011738. N. Prokof’ev and B. Svistunov acknowledge support from the Simons collaboration on Frontiers in Superconductivity. The Flatiron Institute is a division of the simons foundation.
@@ Appendix A Field theoretical approach to the electron-phonon problem
First, we do not approximate the electron-electron interactions using an exchange-correlation potential, as is commonly done in density functional theory (DFT) calculations[giustino_2017]. This allows our derivation to remain applicable to strongly coupled materials. The only assumption we make is that the ions form a rigid lattice structure, implying that the vibration amplitude is small compared to the lattice spacing.
@@ A.1 Electron-ion Model
In this subsection, we present a concise review of the electron-ion problem from first principles, following a derivation adapted from Ref. [stefanucci_2023]. We consider a solid-state system composed of electrons and ions, where the ions form a rigid lattice with equilibrium positions $\mathbf{R}^{0}=(\mathbf{R}^{0}_{1},\ldots,\mathbf{R}^{0}_{N_{n}})$. Due to quantum and thermal fluctuations, the ions oscillate around their equilibrium positions with small amplitudes $\mathbf{u}=(\mathbf{u}_{1},\ldots,\mathbf{u}_{N_{n}})$. At a given time, the ions are located at coordinates $\mathbf{R}\equiv\mathbf{R}^{0}+\mathbf{u}$. The electrons and ions are governed by the full Hamiltonian:
@@ A.5 Density Functional Perturbation Theory of the Electron-Phonon Interaction
We started by reviewing the derivation of the electron-phonon interaction within the framework of DFT. We followed the spirit of dielectric approach [dielectric1, dielectric2], which is theoretically equivalent to the DFPT as pointed out in Ref.[EPW1]. Starting from the effective potential experienced by an electron, we had the Kohn-Sham potential,
@@ Appendix C Finite Temperature Scaling: Generic BCS
As first illustrated in Ref. [pengcheng], for emergent BCS superconductors and non-superconductors at the Fermi surface, the linear response function follows a universal scaling relation, which we call precursory Cooper flow (PCF):
@@ D.1 Phonon effects on the vertices of Quasi-particles
$\int_{\mathbf{k}\omega}A_{\mathbf{k}\omega}\Pi^{\text{s}}_{\mathbf{k}\omega}W^{\text{ph}}_{\mathbf{k}\omega;\mathbf{k}^{\prime}\omega^{\prime}}=\int_{\mathbf{k}\omega}A_{\mathbf{k}_{\text{F}}\omega}\Pi^{\text{s}}_{\mathbf{k}\omega}W^{\text{ph}}_{\mathbf{k}_{\text{F}}\omega;\mathbf{k}^{\prime}\omega^{\prime}}+O(\eta_{\text{e}})$ (136)
@@ E.1 Spin and Angular Momentum Conventions
The UEG model pertains full translational and rotational symmetries, and the superconducting properties depends on the symmetry channel. In this work, we focused on the s-wave spin singlet superconductivity, although the conventions presented below were generic. The generic discussion of angular momentum decomposition of the PCF equation could be found in Appendix A of Ref. [pengcheng].
@@ E.2.1 Self-energy expansion
The self-energy calculation followed the approach explained in Ref. [hou_feynman_2024]. Here we only briefly illustrated the procedure.
@@
where each term $\Sigma^{(n)}$ represents a group of diagrams that could be evaluated stochastically as explained in Ref.[hou_feynman_2024]. The desired quantities, quasiparticle weight and the effective mass, was then given by the equation

### v2

@@ VI Conventional Superconductors
We apply this framework to revisit e-ph SC in simple metals where properties of conduction electrons can be approximated by UEG. Existing predictions of $T_{\text{c}}$ based on the downfolded ME theory or derived approximations, such as the McMillan or Allen-Dynes formulas, result in order-of-magnitude uncertainties for simple metals with sub-Kelvin $T_{\text{c}}$, depending on the choice of the phenomenological Coulomb pseudopotential [69].
@@ VI.1 Methods
Li(9R) 40 242444Electron-phonon coupling in lithium for 9R and hcp structures [59]. We computed the hcp case, but also adapted results from literature for the 9R case. 0.34444Electron-phonon coupling in lithium for 9R and hcp structures [59]. We computed the hcp case, but also adapted results from literature for the 9R case. 0.34444Electron-phonon coupling in lithium for 9R and hcp structures [59]. We computed the hcp case, but also adapted results from literature for the 9R case. 3.25555Data for $r_{\text{s}}$ and $m_{\text{b}}$ in lithium were adapted from Ref.[87]. 1.75555Data for $r_{\text{s}}$ and $m_{\text{b}}$ in lithium were adapted from Ref.[87]. 0.18 $5\times 10^{-3}$ $4\times 10^{-4}$ 0.35
Li(hcp) 41 243 0.33444Electron-phonon coupling in lithium for 9R and hcp structures [59]. We computed the hcp case, but also adapted results from literature for the 9R case. 0.37 3.19 1.4 0.17 0.03 $4\times 10^{-4}$ 0.64
@@
To obtain band structures and phonon spectra we employed Quantum Espresso (QE) [29, 28, 30] package. In all calculations the optimized norm-conserving Vanderbilt (ONCV) potential [37, 94] and Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [73] were used. Metallic crystals considered in this work form either FCC, or BCC, or HCP Bravais lattices. To investigate pressure effects, the lattice constants (specifically for Al) were determined via a fitted equation of state from Ref. [25]; a detailed discussion is provided in the Appendix F. For the self-consistent-field (SCF) calculation, the energy cutoff was set to 90 Ry, the momentum-grid had $\times 24\times 24, 24×24$$\times 24\times 24$, $24\times 24\times 12$ points for the FCC, BCC, and HCP structures, respectively. The error on the converged total energy was less than 0.001 Ry per atom. Phonons were calculated using the DFPT approach of Ref. [9], which was implemented within the QE package. A coarse $\mathbf{q}$-grid for the FCC, BCC, and HCP lattices contained $\times 6\times 6, 6×6$$\times 6\times 6$, and $6\times 6\times 3$ points, respectively, and the acoustic sum rule was used to eliminate the small imaginary frequency at the $\mathbf{\Gamma}$ point.
To get the e-ph coupling constant $\lambda$, we used the EPW package [32, 78], which evaluates e-ph interaction for Wannier functions, and subsequent Fourier interpolation to obtain the e-ph interaction defined on an arbitrary $\mathbf{k},\mathbf{q}$ grid. Wannier functions were generated by the Wannier90 package [76, 64, 97] using coarse $\mathbf{k}$-grids for the FCC, BCC, and HCP lattices with $\times 12\times 12, 12×12$$\times 12\times 12$, and $12\times 12\times 6$ points, respectively. The Wannier projectors for Li, Na, K, Mg, Al were set to $s$ and $p$ orbitals, and the ones for Zn are $s$, $p$, $d$ orbitals. To get a converged result for $\lambda$, we set a fine $\mathbf{k}$ ($\mathbf{q}$) grid for the FCC, BCC, and HCP lattices with $\times 60\times 60, 60×60$$\times 60\times 60$, and $60\times 60\times 30$ points, respectively.
@@ VI.2 Results
Figure 10: The pressure dependence of the superconducting critical temperature in aluminum. The squares are our theoretical results; the lines are guides to the eye. Experimental data from Levy et al. [58] and Gubser et al. [36] are plotted as diamond and circular markers, respectively.
In Fig. 10, we compare the experimental data for the pressure-dependent $T_{\text{c}}$ in aluminum [58, 36] with the corresponding results of our approach (up to the highest reported pressure). We see that our approach accurately captures the experimental trend, showing a clear decrease in $T_{\text{c}}$ as pressure increases from ambient conditions to $6$ GPa.
Although similar benchmarks using the SCDFT method [79] also show good agreement with the experiment, the treatment of Coulomb repulsion within SCDFT involves uncontrolled approximations with respect to vertex corrections, screening dynamics, and quasiparticle weight. In contrast, our approach is based on a rigorous microscopic foundation and clear understanding of how one should proceed in cases significantly different from UEG. We performed calculations beyond the current experimental pressure limit of $6$ GPa; see the inset of Fig. 10 and Fig. 11. Assuming no structural transitions, we predict that SC in Al vanishes at pressure 60~60GPa. Notably, even at $20$ GPa, $T_{\text{c}}$ is already suppressed below $1$ mK, rendering it undetectable with current experimental techniques.
@@
With the same numeric protocol, we calculated $T_{c}$ for other simple metals; the results are presented in Fig. 11 and TABLE 2. Given $T_{c}/\omega_{\text{log}}\propto\exp(-1/g)$ scaling, we choose to plot all data using $y=1/\ln(\omega_{\text{log}}/T\text{c})$ for the vertical axis, which may be considered as an effective attractive coupling in the Cooper channel (if $T_{\text{c}}$ is finite) for a given $\mu^{*}$ and $\lambda$ (here $\omega_{\text{log}}$ is the $\log$-averaged phonon frequency calculated following Ref. [4]).
@@
Our results suggest that Mg, Na, and K are close to the critical point between states with and without s-wave SC. (We do not use the term “quantum critical point” here because SC in some other symmetry channel will preempt it.) Although these metals remain in the normal state as the temperature is lowered (e.g., well below $1K$), they enter a quantum critical regime characterized by a diverging pair-field susceptibility $\chi\sim\ln(T)$ [44] without any fine-tuning. This signal is worth verifying in future experiments.
@@ VII Conclusion
We developed a rigorous ab initio framework, Fig. 9, to obtain downfolded ME theory on the Fermi surface for correlated electrons, which resolves long-standing ambiguities in the treatment of Coulomb repulsion. By systematically integrating out high-energy electronic degrees of freedom using field-theoretical renormalization techniques, we arrive at frequency-only ME equations with all parameters precisely linked to the underlying microscopic Hamiltonian: the Coulomb pseudopotential $\mu^{*}$, determined by the two-quasiparticle scattering amplitude renormalized by the Bogoliubov-Tolmachev-Shirkov logarithm [13, 103], the electron-phonon coupling $\lambda$ screened by the quasiparticle vertex correction, and the quasiparticle density of states. Our approach eliminates the need for phenomenological—and thus uncontrolled—treatment of key parameters and establishes protocols for computing $\mu^{*}$ and $\lambda$ from the first principles.
@@
The ab initio framework of Fig. 9 applies not only to simple metals but to any other correlated material with well-defined quasiparticles at the Fermi surface and large separation between the phonon and electron energy scales $\omega_{\text{D}}\ll E_{\text{F}},\omega_{\text{p}}$. However, implementing high-order diagrammatics and renormalization, as it was done for UEG, is a major technical challenge for lattice systems with complex band structure lacking spherical symmetry. For example, in transition metals more than one band is involved in screening [27] and Umklapp processes cannot be neglected. In cases including two dimensional systems the particle-particle channel Coulomb interaction is non-local (not well approximated by a structureless interaction and as noted by Simonato, Katsnelson and Rosner [96] a more careful treatment is required. If the energy-scale separation condition, on which the downfolding approximation rests, is not satisfied, e.g. at extreme densities relevant to astrophysical objects such as white and black dwarf stars [31], or perhaps in flat-band systems and 2D materials where the $\sqrt{q}$ plasmon dispersion creates challenges to the energy scale separation arguments there may be no real alternative to working with full BSE.
Our findings highlight the crucial role of e-e correlations in the two-particle vertex function; they cannot be quantified by the single-electron exchange-correlation potential within the DFT approximation [41, 53]. Modern numerical methods like VDiagMC enable accurate computation of these correlations, providing key insights into the topic of paring despite strong repulsive forces, and opening new possibilities for understanding other emergent phenomena, such as accurate calculation of many-body forces in metals, with direct implications for more reliable molecular dynamics simulations. In conclusion, by providing a complete and predictive first-principles framework for incorporating e-e correlations into the existing machinery of predicting superconductivity from e-ph interaction, this work represents a step forward in our ability to predict superconductivity starting from properties of the microscopic Hamiltonian without uncontrolled approximations.
@@ Acknowledgements.
We thank S. Das Sarma and A. Chubukov for valuable discussions. K. Chen and X. Cai are supported by the National Natural Science Foundation of China under Grants No. 12474245 and No. 12447103, the National Key Research and Development Program of China under Grant No.2024YFA1408604, and the GHfund A (202407010637). X. Cai and T. Wang acknowledge support from the National Science Foundation under Grant No. DMR-2335904. T. Zhang and S. Zhang acknowledge the support from National Key R&D Project (Grant Nos. 2023YFA1407400 and 2024YFA1409200), and the National Natural Science Foundation of China (Grant Nos. 12374165 and 12447101). A. Millis’s work at Columbia University was supported by the Keele Foundation and the Columbia University Materials Research Science and Engineering Center (MRSEC) through NSF Grant DMR-2011738. N. Prokof’ev and B. Svistunov acknowledge support from the Simons Foundation (SFI-MPS-NFS-00006741-07, N.P. a B.S.) in the Simons Collaboration on New Frontiers in Superconductivity. The Flatiron Institute is a division of the simons foundation.
@@ Appendix A Field theoretical approach to the electron-phonon problem
First, we do not approximate the electron-electron interactions using an exchange-correlation potential, as is commonly done in density functional theory (DFT) calculations[33]. This allows our derivation to remain applicable to strongly coupled materials. The only assumption we make is that the ions form a rigid lattice structure, implying that the vibration amplitude is small compared to the lattice spacing.
@@ A.1 Electron-ion Model
In this subsection, we present a concise review of the electron-ion problem from first principles, following a derivation adapted from Ref. [98]. We consider a solid-state system composed of electrons and ions, where the ions form a rigid lattice with equilibrium positions $\mathbf{R}^{0}=(\mathbf{R}^{0}_{1},\ldots,\mathbf{R}^{0}_{N_{n}})$. Due to quantum and thermal fluctuations, the ions oscillate around their equilibrium positions with small amplitudes $\mathbf{u}=(\mathbf{u}_{1},\ldots,\mathbf{u}_{N_{n}})$. At a given time, the ions are located at coordinates $\mathbf{R}\equiv\mathbf{R}^{0}+\mathbf{u}$. The electrons and ions are governed by the full Hamiltonian:
@@ A.5 Density Functional Perturbation Theory of the Electron-Phonon Interaction
We started by reviewing the derivation of the electron-phonon interaction within the framework of DFT. We followed the spirit of dielectric approach [75, 82], which is theoretically equivalent to the DFPT as pointed out in Ref.[32]. Starting from the effective potential experienced by an electron, we had the Kohn-Sham potential,
@@ Appendix C Finite Temperature Scaling: Generic BCS
As first illustrated in Ref. [44], for emergent BCS superconductors and non-superconductors at the Fermi surface, the linear response function follows a universal scaling relation, which we call precursory Cooper flow (PCF):
@@ D.1 Phonon effects on the vertices of Quasi-particles
$\int_{\mathbf{k}\omega}A_{\mathbf{k}\omega}\Pi^{\text{s}}_{\mathbf{k}\omega}W^{\text{ph}}_{\mathbf{k}\omega;\mathbf{k}^{\prime}\omega^{\prime}}=\int_{\mathbf{k}\omega}A_{\mathbf{k}_{\text{F}}\omega}\Pi^{\text{s}}_{\mathbf{k}\omega}W^{\text{ph}}_{\mathbf{k}_{\text{F}}\omega;\mathbf{k}^{\prime}\omega^{\prime}}+O(\eta_{\text{e}}).$ (136)
@@ E.1 Spin and Angular Momentum Conventions
The UEG model pertains full translational and rotational symmetries, and the superconducting properties depends on the symmetry channel. In this work, we focused on the s-wave spin singlet superconductivity, although the conventions presented below were generic. The generic discussion of angular momentum decomposition of the PCF equation could be found in Appendix A of Ref. [44].
@@ E.2.1 Self-energy expansion
The self-energy calculation followed the approach explained in Ref. [45]. Here we only briefly illustrated the procedure.
@@
where each term $\Sigma^{(n)}$ represents a group of diagrams that could be evaluated stochastically as explained in Ref.[45]. The desired quantities, quasiparticle weight and the effective mass, was then given by the equation